A theater sells children’s tickets for half the adult ticket price. If 5 adult
tickets and 8 children’s tickets cost a total of $27, what is the cost of
an adult ticket?

Answer by Paul(988)

Childern =1/2x , Adult x
x(8)/2+5x=27
4x+5x=27
9x=27
x=3
Hence the cost of one adult ticket is $3.